questionking:

Maya is showing her work in simplifying (5.2 − 4.5) − 5.5 + 4.8. Identify any error in her work or reasoning.

Ultrilliam:

Welcome to QC, love the username BTW :P @Zowie

Zowie:

What are the steps she made? Is this the whole question? I need to whole question.

questionking:

Step 1: (5.2 − 4.5) − 5.5 + 4.8 Step 2: 5.2 + (− 4.5 − 5.5) + 4.8 (identity property) Step 3: 5.2 − 10 + 4.8 Step 4: 5.2 + 4.8 − 10 (commutative property) Step 5: 10 − 10 = 0

questionking:

which one is incorrect

Zowie:

Well you always have to use and follow PEMDAS. So it should look like this.

questionking:

i may need to reset my acount

questionking:

I

Zowie:

(5.2 − 4.5) − 5.5 + 4.8 You do what's in parenthesis first. (0.7)-5.5 + 4.8 Next is exponents. Well in this case there is no exponents. Same as multiplication and division. Next you would do addition. 0.7-10.3 Now the final step is subtraction. -9.6 That is what your answer should be. ^.^

Ultrilliam:

Why is that? Can you not verify your email?

Ultrilliam:

If you can't find the verification email, make sure to check spam!

questionking:

Ok but where did she go wrong?

Zowie:

She went wrong in Step 2.

questionking:

THANK YOU :)

Zowie:

Anytime! ^.^

questionking:

;)

questionking:

TTYGL

Vocaloid:

Nope Mays original work was correct

questionking:

?

Vocaloid:

-5.5 + 4.8 is not -10.3

Zowie:

I didn't say plus

Vocaloid:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Zowie (5.2 − 4.5) − 5.5 + 4.8 You do what's in parenthesis first. (0.7)-5.5 + 4.8 Next is exponents. Well in this case there is no exponents. Same as multiplication and division. Next you would do addition. 0.7-10.3 Now the final step is subtraction. -9.6 That is what your answer should be. ^.^ \(\color{#0cbb34}{\text{End of Quote}}\)

questionking:

not the numbers this part

questionking:

(identity property)

Zowie:

(5.2 − 4.5) − 5.5 + 4.8 (0.7)-5.5 + 4.8 0.7-10.3 -9.6

questionking:

part

Zowie:

That's what it should look like. If you are following the correct way to do math. PEMDAS.

Ultrilliam:

*grabs popcorn*

questionking:

me too

questionking:

:)

Zowie:

When you

Ultrilliam:

*passes popcorn to question*

Vocaloid:

(0.7)-5.5 + 4.8 0.7 - 5.5 + 4.8 = ?

questionking:

guys not the number part

Vocaloid:

1 attachment
questionking:

the (identity) part

Zowie:

When you add together 5.5 and 4.8 you get 10.3. Just because the number is positive it doesn't change the fact that there is a subtraction symbol after 0.7, so you will subtract 10.3 from 0.7

questionking:

what should that part be

Vocaloid:

Go from left to right

Zowie:

Voca Google isn't always right. Google doesn't use PEMDAS. It just goes down the line. When doing math you have to use PEMDAS. No matter what. Question she still messed up on Step 2.

questionking:

no

Zowie:

PEMDAS is the number one rule in math.

questionking:

not the steps

Vocaloid:

I know what pemdas is Addition and subtraction have the same priority under pemdas

Vocaloid:

Look up "left to right rule"

questionking:

Ok never mind I got it

questionking:

chill everyone

questionking:

Which of the following correctly uses absolute value to show the distance between −60 and 13? |−60 − 13| = |−73| = 73 units |−60 − 13| = |−73| = −73 units |−60 + 13| = |−47| = 47 units |−60 + 13| = |−47| = −47 units

questionking:

can you answer this

Vocaloid:

Take the absolute value of the difference

Vocaloid:

-60 - 13 = ?

Vocaloid:

Still there?

Newbreed:

yeah

Newbreed:

I got a new name

Ultrilliam:

I noticed

Vocaloid:

What is -60 - 13 = ?

Newbreed:

-73

Vocaloid:

Good Now we take the absolute value So 73 is the distance

Vocaloid:

So A

Newbreed:

TY

Newbreed:

Which fraction is equivalent to 2 over negative 6?

Vocaloid:

Divide the numerator and denominator by 2

Newbreed:

GTG

Tigerknight:

Woah, that was intense.

Ultrilliam:

Indeed

Vocaloid:

Why do people interpret "Vocaloid is correcting someone's math" as "Vocaloid is angry?" They're just numbers, I'm not emotionally attached to them.

Ultrilliam:

Who did?